Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 across both periods. Let MC = 0. Let reserves = 200. Consider the basic two-period model. What is consumption of the resource in the present?

92.86

100

107.14

115.12

none of the above
Answer #1

Answer option 3) 107.14

Q= 150-P+100

Q= 250-P

using optimal extraction rule of a non renewable resource

250-Q1= (250-Q2)/(1+r)

r = .1

As from resource constraint: Q1+Q2= 200

put Q2= 200-Q1

then , Q1*= 107.143